### v1

@@ VI Conclusion
It should be noted that there are challenges in achieving the sensitivity required to detect GWs, as mentioned in the previous research [19]. However, in the near future, a neutron GW detector will be feasible with the development of neutron optics technology. Additionally, although the current neutron interferometry has not yet reached the level of laser-based systems and is still under development for GW detection, the data analysis methods introduced in this paper enable us to conduct the experiments for a neutron GW detector without waiting for the advancement of neutron optics technology.
@@ VII Acknowledgments
Appendix: Fitting process for experimental results
@@
$+P_{\mathrm{AC}}\sin\bigl(2\pi f_{\mathrm{AC}}t+\phi_{\mathrm{initial}}\bigr{missing})\Bigr{)}\lambda\Bigr{\}}+B_{\mathrm{fixed}}.$ (19)
where $I_{0}(\lambda)$ is used for sample-out measurements to determine $P_{L}$ and $P_{\mathrm{IFO}}$, $I_{1}(\lambda)$ is used for sample-in measurements without modulation to get $P_{\mathrm{sample}}$, and $I_{2}(\lambda)$ is used for sample-in measurements with modulation to obtain $P_{\mathrm{offset}}$, $P_{\mathrm{AC}}$ and $f_{\mathrm{AC}}$. Before each frequency measurement, the yaw angle of the BSE, which maximizes the visibility of the interference fringe, is determined by the yaw-scan measurement. The alignment of the stage position can change the yaw angle of the BSE. The results of yaw-scan measurements are shown in Figs. Appendix: Fitting process for experimental results and Appendix: Fitting process for experimental results.
@@
figureInterference fringes from yaw-scan measurements. In these panels, the red markers represent the data points with error bars, and the black curves are the fitting curves. Each measurement corresponds to a specific position of the yaw stage at the BSE, as shown in each panel.
[…]
figureFitting results from yaw-scan measurements. In both panels, the black lines represent the data points with error bars, and the red lines indicate the fitting curves. The upper and lower panels show the visibility $A_{1}$, the offset $B_{1}$ and the corresponding fitting curves when a step of 1 $\mu$m changes the yaw stage position. A Gaussian curve fits the visibility, and a constant fits the offset.
@@
This setup had already been adjusted to maintain the mid-fringes before the sample-in measurements, based on the yaw-scan measurements. The fitting results for $P_{L}$ and $P_{\mathrm{IFO}}$ from these sample-out measurements are shown in Fig. Appendix: Fitting process for experimental results.
[…]
figureFitting results of $P_{L}$ and $P_{\mathrm{IFO}}$. The black markers represent data points with error bars, while the red lines show the fitting curves. The horizontal axis corresponds to the measurement order. In the left panel, the fitting function is a linear. In the right panel, the fitting function is a constant.
These results show $P_{L}=(0.0027\pm 0.0041)m+(3.231\pm 0.052)$ $\mathrm{nm}$, where $m$ is the measurement number ($m=$1, 2, …, 8), and $P_{\mathrm{IFO}}=-10.71\pm 0.045$ $\mathrm{nm}^{-1}$. $P_{L}$ is fitted by a linear function with the number of sample-in measurements, $m$, because the drift of the geometric optical length is taken into account. In the subsequent fitting processes for the 250 Hz modulation, these $P_{L}$ and $P_{\mathrm{IFO}}$ values are fixed in Eq. (16). Next, $P_{\mathrm{sample}}$ is fitted using a sample-in measurement without modulation. Each sample-in measurement is measured immediately before the modulation of the initial phase to obtain the offset effect from a sample. These results are shown in Figs. 30-31.
@@
These results show $P_{\mathrm{sample}}=15.05\pm 0.012$ $\mathrm{nm}^{-1}$. Since $P_{\mathrm{sample}}$ is a constant during a series of the sample-in measurements, $P_{\mathrm{sample}}$ is fitted as a constant. This $P_{\mathrm{sample}}$ is also fixed in Eq. (16) for the modulation measurements. Finally, the remaining parameters within the cosine function are fitted using the sample-in measurement with modulation, as shown in Fig. Appendix: Fitting process for experimental results.
[…]
figureInterference fringes from the sample-in measurements with the 250 Hz modulation. In these panels, the red markers represent data points with error bars, while the black curves show the fitting results. Each measurement corresponds to a specific initial modulation phase, as shown in each panel.
In this fitting process for the AC parameters, each initial phase $\phi_{\mathrm{initial}}$ is fixed at 45-degree intervals. This approach assumes that the initial phase error is negligible compared to the frequency error. The values for $P_{\mathrm{offset}}$, $P_{\mathrm{AC}}$ and $f_{\mathrm{AC}}$ values obtained from these sample-in measurements are shown in Figs. 32-Appendix: Fitting process for experimental results.
@@
[…]
figureFitting results of $P_{\mathrm{AC}}$ in the left panel and $f_{\mathrm{AC}}$ in the right panel. The black markers represent data points with error bars, and the red lines show the fitting curves.

### v3

@@ VI Conclusion
It should be noted that there are challenges in achieving the sensitivity required to detect GWs, as mentioned in the previous research [19]. While a practical neutron GW detector remains far beyond the reach of current technology, future advances in neutron optics and sources may eventually make such a detector feasible. Importantly, the data analysis methods introduced in this paper enable us to conduct experiments for a neutron GW detector without waiting for advances in neutron optics technology, even though current neutron interferometry has not yet reached the level of laser-based systems and is still under development for GW detection.
@@ VII Acknowledgments
Appendix A Fitting process for experimental results
@@
$+P_{\mathrm{AC}}\sin\bigl(2\pi f_{\mathrm{AC}}t+\phi_{\mathrm{initial}}\bigr.)\Bigr)\lambda\Bigr\}+B_{\mathrm{fixed}}.$ (19)
where $I_{0}(\lambda)$ is used for sample-out measurements to determine $P_{L}$ and $P_{\mathrm{IFO}}$, $I_{1}(\lambda)$ is used for sample-in measurements without modulation to get $P_{\mathrm{sample}}$, and $I_{2}(\lambda)$ is used for sample-in measurements with modulation to obtain $P_{\mathrm{offset}}$, $P_{\mathrm{AC}}$ and $f_{\mathrm{AC}}$. Before each frequency measurement, the yaw angle of the BSE, which maximizes the visibility of the interference fringe, is determined by the yaw-scan measurement. The alignment of the stage position can change the yaw angle of the BSE. The results of yaw-scan measurements are shown in Figs. Appendix A and Appendix A.
Figure 33: Interference fringes from yaw-scan measurements. In these panels, the red markers represent the data points with error bars, and the black curves are the fitting curves. Each measurement corresponds to a specific position of the yaw stage at the BSE, as shown in each panel.
@@
This setup had already been adjusted to maintain the mid-fringes before the sample-in measurements, based on the yaw-scan measurements. The fitting results for $P_{L}$ and $P_{\mathrm{IFO}}$ from these sample-out measurements are shown in Fig. 36.
Figure 36: Fitting results of $P_{L}$ and $P_{\mathrm{IFO}}$. The black markers represent data points with error bars, while the red lines show the fitting curves. The horizontal axis corresponds to the measurement order. In the top panel, the fitting function is a linear. In the bottom panel, the fitting function is a constant.
These results show $P_{L}=(0.0027\pm 0.0041)m+(3.231\pm 0.052)$ $\mathrm{nm}$, where $m$ is the measurement number (m=m=1, 2, …, 8), and $P_{\mathrm{IFO}}=-10.71\pm 0.045$ $\mathrm{nm}^{-1}$. $P_{L}$ is fitted by a linear function with the number of sample-in measurements, $m$, because the drift of the geometric optical length is taken into account. In the subsequent fitting processes for the 250 Hz modulation, these $P_{L}$ and $P_{\mathrm{IFO}}$ values are fixed in Eq. (16). Next, $P_{\mathrm{sample}}$ is fitted using a sample-in measurement without modulation. Each sample-in measurement is measured immediately before the modulation of the initial phase to obtain the offset effect from a sample. These results are shown in Figs. 37-38.
@@
These results show $P_{\mathrm{sample}}=15.05\pm 0.012$ $\mathrm{nm}^{-1}$. Since $P_{\mathrm{sample}}$ is a constant during a series of the sample-in measurements, $P_{\mathrm{sample}}$ is fitted as a constant. This $P_{\mathrm{sample}}$ is also fixed in Eq. (16) for the modulation measurements. Finally, the remaining parameters within the cosine function are fitted using the sample-in measurement with modulation, as shown in Fig. Appendix A.
Figure 39: Interference fringes from the sample-in measurements with the 250 Hz modulation. In these panels, the red markers represent data points with error bars, while the black curves show the fitting results. Each measurement corresponds to a specific initial modulation phase, as shown in each panel.
In this fitting process for the AC parameters, each initial phase $\phi_{\mathrm{initial}}$ is fixed at 45-degree intervals. This approach assumes that the initial phase error is negligible compared to the frequency error. The values for $P_{\mathrm{offset}}$, $P_{\mathrm{AC}}$ and $f_{\mathrm{AC}}$ values obtained from these sample-in measurements are shown in Figs. 40-41.
@@
Figure 41: Fitting results of $P_{\mathrm{AC}}$ in the top panel and $f_{\mathrm{AC}}$ in the bottom panel. The black markers represent data points with error bars, and the red lines show the fitting curves.